A modern-day zeppelin holds 9,130 m3 of helium. Compute its maximum payload at sea level. (Assume the helium and air to be at 0°C and 1 atm.)


1
Expert's answer
2020-07-28T10:27:30-0400

For this problem we are given the volume in cubic meters, therefore we will use the fact that the weight density of a substance can be represented as the mass density of the substance times the acceleration due to gravity:


"Dw = Dg"


Step 1:

Volume:

V = 9130 "m^{3}"

Mass Density of air:

D(air) = 1.29 "( \tfrac{m}{m^{3}})"

Mass Density of helium:

D(helium) = 0.18 "( \tfrac{m}{m^{3}})"


Step 2:

Weight of helium:

W = ?

Buoyant force:

F(buoyant) = ?

Maximum payload or Net Force:

Fnet = ?


Step 3:

Weight of helium:

"W = D_{helium}\cdot g \cdot V_{helium}"

Buoyant force:

"F_{buoyant} = D_{air} \cdot g \cdot V_{air}"


In this case "V_{helium} = V_{air} = V"


Net Force:

"F_{net} = F_{buoyant} - W"


Step 4:

Net Force:


"F_{net} = F_{buoyant} - W=D_{air} \cdot g \cdot V-D_{helium}\cdot g \cdot V=[D_{air} -D_{helium}]\cdot g \cdot V"


Step 5:

Final Answer:


"F_{net} =" "(1.29 - 0.18) \cdot g \cdot 9130 = 99,417.4 N"


"F_{net} =" "99,417.4 N"